# citation.bib

### v1

@@ Abstract
The rapid advancement of large language models (LLMs) necessitates the development of new benchmarks to accurately assess their capabilities. To address this need for Vietnamese, this work aims to introduce ViLLM-Eval, the comprehensive evaluation suite designed to measure the advanced knowledge and reasoning abilities of foundation models within a Vietnamese context. ViLLM-Eval consists of multiple-choice questions and predict next word tasks spanning various difficulty levels and diverse disciplines, ranging from humanities to science and engineering. A thorough evaluation of the most advanced LLMs on ViLLM-Eval revealed that even the best performing models have significant room for improvement in understanding and responding to Vietnamese language tasks. ViLLM-Eval is believed to be instrumental in identifying key strengths and weaknesses of foundation models, ultimately promoting their development and enhancing their performance for Vietnamese users.  [^1] The ViLLM-Eval data and evaluation code are available at https://huggingface.co/datasets/vlsp-2023-vllm/lm-evaluation-harness.
@@ 1 Introduction
To address this gap, this work introduces ViLLM-Eval, a comprehensive evaluation suite meticulously designed to thoroughly assess the advanced knowledge and reasoning abilities of foundation models within the Vietnamese context. ViLLM-Eval comprises multiple-choice questions and predict-next-word tasks spanning various difficulty levels and diverse disciplines, ranging from humanities to science and engineering. The questions have been carefully curated to reflect the knowledge and reasoning skills relevant to Vietnamese users, including aspects of Vietnamese culture, history, and current affairs.
@@ 3.1 Design Principle
Name dataset Num samples
LAMBADA 10,246
Exam 19,150
Gen. Knowl. 2,000
Comp. QA 900
@@ 6 Discussion
In conclusion, while ViLLM-Eval represents a significant step forward in evaluating LLMs within the Vietnamese language context, there remains ample room for further research and development. By addressing the limitations of current evaluation approaches and embracing a more holistic perspective, this work aims to unlock the full potential of LLMs and pave the way for their responsible and effective deployment in real-world applications.
@@ 8 Acknowledgements
This work is supported by Intelligent Integration Co., Ltd (INT2), Ha Noi, Vietnam and HPC SYSTEMS Inc., Tokyo, Japan. We also would like to express our sincere gratitude to the students: Nguyen Trong Chi, Nguyen Phuoc Nguyen, Vo Huu Tri and Nguyen Minh Trung Hieu from the NLP-KD Lab, Ton Duc Thang University. They have contributed to the completion of this research.

### v2

@@ Abstract
The rapid advancement of large language models (LLMs) necessitates the development of new benchmarks to accurately assess their capabilities. To address this need for Vietnamese, this work aims to introduce ViLLM-Eval, the comprehensive evaluation suite designed to measure the advanced knowledge and reasoning abilities of foundation models within a Vietnamese context. ViLLM-Eval consists of multiple-choice questions and predict next word tasks spanning various difficulty levels and diverse disciplines, ranging from humanities to science and engineering. A thorough evaluation of the most advanced LLMs on ViLLM-Eval revealed that even the best performing models have significant room for improvement in understanding and responding to Vietnamese language tasks. ViLLM-Eval is believed to be instrumental in identifying key strengths and weaknesses of foundation models, ultimately promoting their development and enhancing their performance for Vietnamese users [^1] The ViLLM-Eval data and evaluation code are available at https://huggingface.co/datasets/vlsp-2023-vllm/ViLLM-Eval. This paper provides a thorough overview of ViLLM-Eval as part of the Vietnamese Large Language Model shared task, held within the 10th International Workshop on Vietnamese Language and Speech Processing (VLSP 2023) [^2] https://vlsp.org.vn/vlsp2023.
@@ 1 Introduction
In this paper, we introduce ViLLM-Eval, a comprehensive evaluation suite meticulously designed to thoroughly assess the advanced knowledge and reasoning abilities of foundation models within the Vietnamese context, as part of the Vietnamese Large Language Models challenge at VLSP 2023. ViLLM-Eval comprises multiple-choice questions and predict-next-word tasks spanning various difficulty levels and diverse disciplines, ranging from humanities to science and engineering. The questions have been carefully curated to reflect the knowledge and reasoning skills relevant to Vietnamese users, including aspects of Vietnamese culture, history, and current affairs.
@@ 3.1 Design Principle
Name dataset Num samples Description
LAMBADA_vi 10,246 The LAMBADA dataset evaluates LLMs’ contextual reasoning ability by challenging them to predict the final word of a sentence within a paragraph consisting of 4-6 sentences.
Exam 19,150 The Exam Vietnamese dataset includes multiple-choice questions spanning various subjects from grades 6 to 12, providing insight into the model’s academic proficiency across diverse domains.
Gen. Knowl. 2,000 The General Knowledge dataset comprises multiple-choice questions derived from popular Vietnamese television shows, testing the model’s breadth of worldly awareness.
Comp. QA 900 The Comprehension QA dataset simulates an in-depth reading comprehension scenario, requiring the model to parse and understand extensive narratives before selecting the correct answer.
@@ 6 Discussion
In conclusion, the VLSP 2023 - Vietnamese Large Language Models shared task was formulated to propel research development in the domain of Large Language Models in Vietnamese. By addressing the limitations of current evaluation approaches and embracing a more holistic perspective, this work aims to unlock the full potential of LLMs and pave the way for their responsible and effective deployment in real-world applications.
@@ 8 Acknowledgements
This work is supported by Intelligent Integration Co., Ltd (INT2), Ha Noi, Vietnam and HPC SYSTEMS Inc., Tokyo, Japan. We would like to express to the organizers of the Vietnamese Language and Speech Processing (VLSP) for their exceptional dedication in orchestrating the Vietnamese Large Language Models challenge. Their efforts have been instrumental in fostering advancements in the field. We also would like to express our sincere gratitude to the students: Nguyen Trong Chi, Nguyen Phuoc Nguyen, Vo Huu Tri and Nguyen Minh Trung Hieu from the NLP-KD Lab, Ton Duc Thang University. They have contributed to the completion of this research.